### v3

# This version: )

Abstract
We consider the standard moral hazard problem with limited liability. The first-order approach (FOA) is the main tool for its solution, but existing sufficient conditions for its validity are restrictive. Our main result shows that the FOA is broadly valid, as long as the agent’s reservation utility is sufficiently high. In basic examples, the FOA is valid for almost any positive reservation wage.
We establish existence and uniqueness of the optimal contract. We derive closed-form solutions with various functional forms. We show that optimal contracts are either linear or piecewise linear option contracts with log utility and output distributions in an exponential family with linear sufficient statistic (including Gaussian, exponential, binomial, geometric, and Gamma). We provide an algorithm for finding the optimal contracts both in the case where the FOA is valid and in the case where it is not at trivial computational cost.
1 Introduction
One of the workhorse models in economics is the principal-agent problem with moral hazard. The principal hires the agent to take an action $a$ in $\mathbb{R}_{+}$ that affects the distribution $f(y|a)$ of output $y$. The principal can only condition payments on realized output, and designs a contract $w(y)\geq 0$ to provide incentives to the agent. The agent chooses the action $a$ to maximize her utility from wages minus her cost of effort $c(a)$ and the principal chooses the contract to maximize expected profits.
The main solution method for this problem is the first-order approach, which assumes that only local deviations in $a$ are binding, and yields a simple formula for the optimal contract. [^1] See [20], the excellent survey [16], and recent work in [14, 27, 6, 8, 4]. Because of tractability, a vast literature simply assumes that the first-order approach is valid, and many applied papers make restrictive assumptions to avoid the issue of non-local deviations.222[29] showed that the first-order approach is not always valid. Examples of papers assuming the first-order approach include [23, 30, 8, 4]. Virtually all early work, including the seminal papers by [20, 21], and [38] assumes it. [21] notes that “one has to assume that the agent’s optimal choice of action is unique for the optimal […] This assumption seems very difficult to validate […] and regrettably we have to leave the question about its validity open.” In applied work, restrictive assumptions such as linear contracts or binary effort are often made to avoid this issue. [15] is an example of an important recent paper using both binary effort and linear contracts.
Unfortunately, existing sufficient conditions for the first-order approach to be valid are restrictive. The seminal papers are [35] and [24], followed by an extensive literature. [^3] Important generalizations include [36, 14, 25, 6], and [26]. [8, 9] develop sufficient conditions with limited liability. [27] summarize the general view that “conditions facilitating the first-order approach are typically quite demanding.” The key issues are elegantly explained by [8] and [14], who says “Unfortunately, the Jewitt conditions are tied to the concavity, not only of the technology, but also of the payment schedule. Thus, there are interesting cases where the Jewitt conditions should fail, since the payment schedule is often not concave. For example, managers often receive stock options, face liquidity constraints.”444[14] continues: “ I believe that it is natural to expect the first-order approach itself to fail in such cases, since the agent’s overall objective function will tend to be nonconcave […]. The fact that CDFC and CISP imply concavity of the agent’s payoff, regardless of the curvature of [the payment], then suggests that the CDFC/CISP conditions are very restrictive, not that the first-order approach is widely applicable.”
Our main result, Theorem 1, shows that the first-order approach is broadly valid as long as the agent’s reservation utility is sufficiently high. The first-order approach holds even when contracts are option-like and the agent’s problem has multiple local maxima. Theorem 1 also implies that the optimal contract exists, is unique, and is characterized by the standard simple formula. The main substantial condition is that the score $\partial_{a}\log f(y|a)$ of the output distribution is increasing in output.
The basic idea is simple, and clearly illustrated in basic examples as in Figure 1. When reservation utility is very low, the first-order approach often fails. Optimal contracts sometimes pay zero unless effort is quite high. This leads agents to be indifferent between equilibrium effort and giving up to nearly zero effort. However, as reservation utility increases, payoffs close to the intended action become more attractive, and the first-order approach becomes valid. Formally, the proof of Theorem 1 is built around the kink in the relaxed optimal contract moving far to the left, and this making the agent’s problem more concave.
Figure 1: Optimal contracts with Gaussian distribution and log utility.
Note: Top panel: optimal wage function $w(y)$. Bottom panel: agent’s expected utility $U(v^{*},a)$ given optimal contract and action $a$. Colors represent reservation utility. Dashed lines indicate that the first-order approach is invalid at that reservation utility. Dots indicate the recommended action. The thin horizontal line indicates indifference between local maxima. Output has gaussian distribution with mean $a$ and standard deviation 5050, initial wealth is 5050 (both in thousands of dollars), and the cost function is $c(a)=a^{2}/30000$.
Section 2 explains our results through simple examples and reconciles our findings with the literature. Section 3 gives definitions. Section 4 states Theorem 1. Section 5 outlines the proof. Section 6.1 provides a broad range of closed-form solutions for optimal contracts. Section 6.2 shows that optimal contracts are piecewise linear option contracts for log utility and output distributions in an exponential family with linear sufficient statistic (this has been independently discovered by [31]). Section 6.3 provides an algorithm for finding the optimal contracts. Section 6.4 gives counter-examples and discusses limitations of our results. Section 6.5 discusses the relationship between our results and the literature.
2 Clarifying Example
Consider the following Gaussian-log utility example. A risk-neutral principal hires an agent to perform a task. The agent chooses action $a$ at a cost $c(a)$ proportional to $a^{2}$. Output $y$, which accrues to the principal, is normally distributed with mean $a$ and standard deviation $\sigma$. The agent has some initial wealth and log utility, so that her utility from a payment of $x\geq 0$ is $u(x)=\log(w_{0}+x)$. We consider $a$ choices in the order of \$100,000, standard deviation $\sigma$ of \$50,000, and $w_{0}$ equal to \$50,000. The agent has reservation wage $\bar{w}$.
The principal designs a compensation contract with wage $w(y)\geq 0$ to maximize profits. The optimal contract balances the goals of inducing effort, reducing the agent’s risk, and providing the agent with her reservation utility.
Figure 1 plots optimal contracts for a range of reservation wages $\bar{w}$, with certainty equivalents from \$0 to \$50,000. The top panel plots the optimal wage schedules $w(y)$. The bottom panel plots the agent’s expected utility $U(v^{*},a)$ of an optimal contract $v^{*}$ as a function of action $a$. The intended action is denoted by the red dot in the bottom panel.
The bottom panel shows for what reservation wages the first-order approach is valid. The first-order approach fails whenever there is a binding non-local deviation. In these cases, solving the problem with only the local incentive constraint leads to incorrect solutions. Cases in which the first-order approach fails are denoted with dashed lines.
The first-order approach fails with a reservation wage of $\mathdollar 1,000$, but is valid for all the illustrated positive reservation wages. This is an illustration of our main point, that the first-order approach is broadly valid in most interesting cases. Moreover, this is representative of a more general phenomenon, and not simply this example. Figures 2 and 3 consider a different distribution (the discrete Poisson distribution) and a different utility function (CARA). The same pattern holds, as it does in a multitude of other examples. Readers can explore more examples in our interactive demo: https://eduardomazevedo.github.io/azevedo-wolff-2025-site/.
Figure 2: Optimal contracts with Poisson distribution and log utility.
Note: Top panel: optimal wage function $w(y)$. Bottom panel: agent’s expected utility $U(v^{*},a)$ Colors represent reservation utility. Dashed lines indicate that the first-order approach is invalid at that reservation utility. Bottom panel dots indicate the recommended action. The thin horizontal line indicates indifference between local maxima. Number of successes has Poisson distribution with mean $a$ and output is one thousand dollars per success, initial wealth is 5050 (in thousands of dollars), and the cost function is $c(a)=a^{2}/399$.
Figure 3: Optimal contracts with Gaussian distribution and CARA utility.
Note: Top panel: optimal wage function $w(y)$. Bottom panel: agent’s expected utility $U(v^{*},a)$ Colors represent reservation utility. Dashed lines indicate that the first-order approach is invalid at that reservation utility. Bottom panel dots indicate the recommended action. The thin horizontal line indicates indifference between local maxima. Parameters are as in figure 1, but with CARA coefficient set to relative risk aversion of 11 at the initial wealth.
Indeed, Theorem 1 shows that the first-order approach is valid for sufficiently high reservation utility. The examples illustrate the theorem idea in simple terms. For very low reservation utility, the first-order approach fails. The reason is that principal offers tough contracts, with zero pay for low output. Optimal contracts have a kink, and only start paying a positive wage relatively close to the intended action. Thus, the agent has a global deviation to basically give up and exert no effort. Thus, the first-order approach fails with very low reservation utility.
However, as reservation utility increases, the first-order approach becomes valid. The reason is that the principal is forced to offer contracts with higher utility at the recommended action. This constraint moves the rightmost local maximum up. The rightmost maximum goes up faster than the leftmost local maximum. Thus, the rightmost maximum becomes the unique global maximum. In the examples, the first-order approach is valid for any positive reservation wage. The proof of Theorem 1 works with a somewhat different logic. The key observation is that the kink in the relaxed optimal contract moves to the left. This then implies that the agent’s problem becomes more concave, and the first-order approach becomes valid.
Theorem 1 does not contradict existing results in the literature. The low reservation utility confirms the literature consensus that it is hard to guarantee that the first-order approach is valid is correct – if we insist that the first-order approach is valid for all reservation utilities. We see that, even in the Gaussian-log utility case, the first-order approach fails. Thus, any condition guaranteeing the first-order approach for all reservation utilities rules out basic examples and is therefore quite restrictive. The examples formally demonstrate the point made by [8] and [14], that requiring the first-order approach to be valid for all reservation utilities excludes many interesting cases.
As a corollary to Theorem 1, the problem is often tractable because we can use standard first-order approach arguments dating back to [20]. We show that an optimal contract exists, is unique, and optimal wages have a simple formula:
$w(y)=k\circ g\biggl(\mu+\lambda S(y|a_{0})\biggr)\text{.}$ (1)
The function $k\circ g$ is determined by the utility function and the limited liability constraint. The score function $S(y|a_{0})$ is determined by the distribution of output. The optimal action is $a_{0}$, and $\lambda$ and $\mu$ are Lagrange multipliers. We give closed-form solutions for these functions for a wide range of examples, and use these formulas to develop an algorithm to compute optimal contracts.
Another corollary is that optimal contracts are piecewise linear in a variety of examples. With log utility, $k\circ g$ is piecewise linear. It equals zero up to a point, and then increases linearly. And output distribution in an exponential family with linear sufficient statistic have linear score functions. Thus, with both of these conditions, the optimal contract is piecewise linear. Figures 1 and 2 illustrate this in the Gaussian and Poisson cases. Note that the optimal contract is only piecewise linear when the first-order approach is valid. This is clear in Figure 1, where the optimal contract is noticeably non-linear for reservation wage of $\mathdollar 1,000$, but otherwise piecewise linear. The CARA case in Figure 3 illustrates that, without log utility, the optimal contract is not piecewise linear.
3 Model
3.1 Model
A risk-neutral principal hires an agent with limited liability. The agent has utility u(x)c(a)u(x)-c(a) from receiving a payment $x\in\mathbb{R}_{+}$ and taking action $a$ in $\mathcal{A}\subseteq\mathbb{R}_{+}$. Both $u$ and $c$ are strictly increasing. Output $y\in\mathbb{R}$ depends on the agent’s action and is distributed according to the output density $f(y|a)$.
A contract is a function $v:\mathbb{R}\rightarrow u(\mathbb{R}^{+})$, which specifies the agent’s utility $v(y)$ from the payment as a function of output $y$. Contracts are defined in terms of utility to simplify notation. To define the wage payments, let the compensation cost function $k$ be the inverse of $u$. A wage function is a function $w:\mathbb{R}\rightarrow\mathbb{R}^{+}$. The wage function associated with contract $v$ is $w(y)=k(v(y))$. Let $\mathcal{C}$ be the set of feasible contracts.
The agent’s utility from a contract $(v)$ given action $(a)$ is
$U(v,a):=\int v(y)f(y|a)\,dy-c(a)\text{.}$
The expected wage is
$W(v,a):=\int k(v(y))f(y|a)\,dy\text{.}$
The cost minimization problem, given an intended action $a_{0}\in\mathcal{A}$ and reservation utility $\bar{U}$, is to choose a contract $v$ to minimize the expected wage subject to the individual rationality (IR) and global incentive compatibility (GIC) constraints. Individual rationality requires that the agent’s expected utility is at least the reservation utility, and global incentive compatibility requires that choosing $a_{0}$ is optimal for the agent. The cost minimization problem is to choose $v$ in $\mathcal{C}$ to
minimize $W(v,a_{0})$
subject to $U(v,a_{0})\geq\bar{U},$ (IR)
$U(v,a_{0})\geq U(v,\hat{a})\quad\forall\hat{a}\in\mathcal{A}.$ (GIC)
The relaxed cost minimization problem replaces the global incentive compatibility constraint with the local incentive compatibility constraint (LIC):
minimize $W(v,a_{0})$
subject to $U(v,a_{0})\geq\bar{U},$ (IR)
$\partial_{a}U(v,a_{0})=0.$ (LIC)
3.2 Assumptions and Notation
We now state the needed assumptions. This section can be skimmed on a first reading.
Assumption 1 (Regularity of utility, cost, and density).
The set of feasible actions $\mathcal{A}\subseteq\mathbb{R}_{+}$ is a compact interval including 0. The utility function $u$ is strictly concave and smooth, has $\lim_{x\rightarrow\infty}u(x)=\infty$ and $\lim_{x\rightarrow\infty}u^{\prime}(x)=0$. The cost function $c$ is strictly convex and smooth, and $c^{\prime}(0)=0$. The output density $f(y|a)$ is smooth in the action $a$ for all outputs $y$.
We now make technical assumptions that allow us to use Leibniz’s rule of differentiation under the integral sign. For this, we impose restrictions on $k$, $f$, and on the set of feasible contracts. At the same time, we require the set of feasible contracts to be rich enough so that relevant contracts are not ruled out by assumption.
Assumption 2 (Regularity of feasible contracts).
  1. 1.
    Every feasible contract is measurable.
  2. 2.
    Given any $a_{1}$ in $\mathcal{A}$ and a feasible contract $v$, there exists an integrable function $\theta(y)$ and a neighborhood of $a_{1}$ such that, for all $a$ in this neighborhood,
    $\left|v(y)\partial^{n}_{a}f(y|a)\right|\leq\theta(y)$
    for $n=0,1,2$ and for all $y$ in the support of $f(\cdot|a)$.
  3. 3.
    For all $a$ in $\mathcal{A}$ and canonical contract $v$ (see definition 1 below),
    $W(v,a)<\infty\text{.}$
  4. 4.
    The set of feasible contracts is convex and includes all canonical contracts.
We now consider substantive restrictions on the distribution of output. [^5] The most unusual assumption is the third one. It holds for example in Gaussian, logistic, and Student-$t$ location families. Examples that do not satisfy the assumption can be constructed with families of distributions where this second derivative oscillates in the left tail.
Assumption 3 (Assumptions on the distribution of output).
  1. 1.
    The support of $f(\cdot|a)$ is the real line.
  2. 2.
    For all $a$, the score $\partial_{a}\log f(y|a)$ is strictly increasing in $y$ and its image is $\mathbb{R}$ (this is known in the literature as the monotone likelihood ratio property or MLRP).
  3. 3.
    There exists $y_{0}$ in $\mathbb{R}$ such that, for all $y\leq y_{0}$ and $a$ in $\mathcal{A}$,
    $f_{aa}(y|a)>0\text{.}$
Another substantial restriction is sufficient concavity of $k^{\prime-1}$.
Assumption 4 (Concavity and limit of inverse marginal utility).
The function ${k^{\prime}}^{-1}$ is strictly concave and
$\lim_{z\to\infty}z\frac{d}{dz}{k^{\prime}}^{-1}(z)$
is finite.
[…]
4 Main Result
We say that the first-order approach is valid for an action and reservation utility pair $(a_{0},\bar{U})$ if any solution to the relaxed cost minimization problem is a solution to the cost minimization problem.
Our main result is that the first-order approach is valid for sufficiently high reservation utility:
Theorem 1.
[Validity of the first order approach with high reservation utility] Given an action $a_{0}>0$, the first-order approach is valid for $(a_{0},\bar{U})$ for any sufficiently high reservation utility $\bar{U}$.
The theorem formalizes the main point discussed in the introduction and examples. That the first order approach is valid under reasonable assumptions, as long as reservation utility is high. In particular, the theorem implies that the cost minimization problem has a tractable solution, given by the standard formulae in the first-order approach literature. We establish the solution in Proposition 1 below, and note the implication to the full problem here.
Corollary 1.
[Solution of the cost minimization problem] Given an action $a_{0}>0$, for any sufficiently high reservation utility $\bar{U}$, the cost minimization problem has an almost everywhere unique solution given by the formula in Proposition 1.
5 Proof of the Main Result
The proof of the main result follows from two key propositions. Proposition 1 characterizes the solution of the relaxed problem. Proposition 2 shows that, for sufficiently high reservation utility, the agent’s problem is concave. We now state these propositions and explain the key steps in the argument. Appendix A contains the proofs.
For the relaxed problem to be well-defined, the derivative $\partial_{a}U(v,a)$ must exist. Remark 2 shows that this is true under our assumptions, and moreover, that we can calculate the derivative by differentiating under the integral sign. Henceforth, we will use differentiation under the integral sign without referencing Remark 2.
Throughout this section, fix a regular intended action $a_{0}$. To simplify notation, we omit the dependence on $a_{0}$ whenever it is clear, writing, for example, $U(v)$ instead of $U(v,a_{0})$.
5.1 Solution to the Relaxed Problem
We first show that the relaxed cost minimization problem has an almost everywhere unique solution, and that this solution has a simple formula in terms of Lagrange multipliers. This section is a minor extension of standard results on the relaxed problem going back to [20] and [23].
The relaxed cost minimization problem is convex. Define its Lagrangian as
$\mathcal{L}(v,\lambda,\mu):=W(v)+\lambda\left(\bar{U}-U(v)\right)+\mu(-U_{a}(v))\text{.}$ (2)
Heuristically differentiating this Lagrangian with respect to $v(y)$ and setting the derivative to zero gives
$k^{\prime}(v(y))f(y|a)=\lambda f(y|a)+\mu f_{a}(y|a_{0})\text{.}$
Dividing by $f(y|a)$ gives
$k^{\prime}(v(y))=\lambda+\mu\frac{f_{a}(y|a_{0})}{f(y|a_{0})}\text{.}$ (FOC)
Equation (FOC) is the key step in the standard first-order approach literature. Appendix A formally analyzes the convex program, demonstrates existence and uniqueness of the solution, and characterizes the optimal contract based on equation (FOC).
The solution is best described with the following notation. Define the optimal expected wage $\omega(\bar{U})$ as the value of the infimum in the relaxed cost minimization problem. Define the link function $g:\mathbb{R}\rightarrow\mathbb{R}$ as [^6] The link function’s input, $z$, is a marginal dollar cost of providing one util to the agent (measured in units of $\frac{\text{\textdollar}}{\text{util}}$). The link function evaluated at $z$, $g(z)$, returns the utility level where $z$ is the marginal cost of utility to the agent. Any $z$ below $\frac{1}{u^{\prime}(0)}$, which is the cheapest possible marginal cost, returns $g(z)=u(0)$.
$g(z):=k^{\prime-1}\left(\max\left\{\frac{1}{u^{\prime}(0)},z\right\}\right)\text{.}$
Define the score function [^7] The score is known as the likelihood ratio in the economics literature. We favor the term score to be in line with the modern scientific literature. as in the statistics literature:
$S(y|a):=\frac{\partial}{\partial a}\log f(y|a)=\frac{f_{a}(y|a)}{f(y|a)}\text{.}$
Definition 1.
A canonical contract $V(y|\lambda,\mu)$ is defined for $\lambda$ and $\mu$ in $\mathbb{R}$ as
$V\left(y|\lambda,\mu\right):=g\biggl(\lambda+\mu S(y|a_{0})\biggr)\text{.}$ (3)
The following proposition shows that the relaxed problem has a solution and characterizes the solution and Pareto frontier.
Proposition 1.
[Solution of the Relaxed Cost-Minimization Problem] There exists $\bar{U}_{L}$ in $\mathbb{R}$ such that:
  1. 1.
    (existence, uniqueness, and characterization) The relaxed problem has an almost everywhere unique solution $v^{*}(y|\bar{U})$. There exist Lagrange multipliers $\lambda^{*}(\bar{U})\geq 0$ and $\mu^{*}(\bar{U})>0$ such that the solution is almost everywhere equal to the canonical contract
    $v^{*}(y|\bar{U}):=V\biggl(y\big|\lambda^{*}(\bar{U}),\mu^{*}(\bar{U})\biggr)\text{.}$
  2. 2.
    (comparative statics)
    • For $\bar{U}\leq\bar{U}_{L}$, we have $\lambda^{*}(\bar{U})=0$. The relaxed optimal contract $v^{*}(\cdot,\bar{U})$ and optimal expected wage $\omega(\bar{U})$ do not vary with $\bar{U}$ in this range.
    • For $\bar{U}>\bar{U}_{L}$, $\lambda^{*}(\bar{U})$ is strictly increasing and $\lim_{\bar{U}\rightarrow\infty}\lambda^{*}(\bar{U})=\infty$. The optimal expected wage $\omega(\bar{U})$ is strictly increasing and strictly convex.
The proposition shows that the solution to the relaxed cost minimization problem is unique and given by a canonical contract. Moreover, the Pareto frontier of profits and agent utility is convex.
The Gaussian-log utility example illustrates proposition 1. The score is $(y-a)/\sigma^{2}$, the compensation cost function is $k(v)=\exp(v)-w_{0}$, and the link function is $\log\max\{z,w_{0}\}$. This implies the wage function
$k\left(v^{*}(y|\bar{U})\right)=\left[\lambda^{*}(\bar{U})+\mu^{*}(\bar{U})\frac{y-a}{\sigma^{2}}-w_{0}\right]^{+}\text{.}$
This is the piecewise linear solution that we saw numerically in figure 1. The convex Pareto frontier is illustrated in figure 4.
Figure 4: Pareto frontier in the Gaussian-log utility example
Note: This figure displays the Pareto frontier of the cost minimization problem and the relaxed cost minimization problem in the Gaussian-log utility example. The first-order approach is valid for all reservation utilities where the two sets coincide. Parameters are as in figure 1.
The formula in proposition 1 is simply solving the first-order condition (FOC) accounting for limited liability. The intuition is the following: the principal would like to always pay the agent $g(\lambda^{*}(\bar{U}))$, which is the utility level where the marginal cost of providing utility to the agent is $\lambda^{*}(\bar{U})$. However, incentive compatibility requires that payment depends on whether there is statistical evidence of high effort. The $\mu$ term in the optimal contract pays more if the score is positive and less if the score is negative. Finally, the maximum ensures that limited liability is respected.
5.2 High Reservation Utility
We now demonstrate that, for sufficiently high reservation utility, the solution to the relaxed cost minimization problem also solves the original cost minimization problem. To do so, we show that the relaxed problem’s solution satisfies the global incentive compatibility constraint. This is achieved by demonstrating the stronger result that the agent’s utility $U(v^{*},a)$ is concave in $a$ at the relaxed optimal contract.
Proposition 2.
[Concavity of the Agent’s Problem for High Reservation Utility] There exists $U^{*}$ in $\mathbb{R}$ such that, for all $\bar{U}\geq U^{*}$ and all $a\in\mathcal{A}$,
$U_{aa}(v^{*}(y|\bar{U}),a)\leq 0.$
Theorem 1 is a direct consequence of this fact:
Proof of Theorem 1.
Take $\bar{U}\geq U^{*}$. If the relaxed cost minimization problem has a solution $v^{*}(y|\bar{U})$, proposition 2 implies that the agent’s expected utility $U(v^{*},a)$ is concave in $a$. Therefore, the global incentive compatibility constraint is satisfied, and the solution to the relaxed cost minimization problem is also a solution to the cost minimization problem. ∎
Figure 5 illustrates the proposition in the Gaussian-log utility example. We use the same parameters as in figure 1, but consider the cost minimization problem with intended action $a_{0}=\mathdollar 100,000$. We see the same pattern as in figure 1. For low reservation utility, there are multiple local maxima, and the first-order approach is invalid. For high reservation utility, there is only one local maximum, and the first-order approach is valid.
Figure 5: Optimal contracts with Gaussian distribution and log utility in cost minimization problem with intended action $a_{0}=\mathdollar 100,000$.
Note: Top panel: optimal wage function $w(y)$. Bottom panel: agent’s expected utility $U(v^{*},a)$ given optimal contract and action $a$. Colors represent reservation utility. Dashed lines indicate that the first-order approach is invalid at that reservation utility. Dots indicate the recommended action. The thin horizontal line indicates indifference between local maxima. Intended action is $a_{0}=\mathdollar 100,000$ and other parameters are as in figure 1.
The intuition for why the agent’s problem is concave for high reservation utility is as follows. The relaxed optimal contract has two regions. When the outcome $y$ is below a threshold, the limited liability constraint binds, and the contract specifies a constant wage of 0. When $y$ is above the threshold, the agent’s payment is dictated by equation (FOC). This kink introduces convexity into the agent’s problem. As reservation utility increases, the kink moves to the left, so that the agent is paid with high probability. In the example, $a_{0}$ is fixed at \$100,000. As the kink moves to the left, the probability that the agent is paid a non-zero wage conditional on $a_{0}$ approaches $1$.
Appendix A formalizes this argument in the proof of proposition 2. The proof has two key steps. First, it is shown that the probability that the agent receives zero payment converges to zero. When $\bar{U}$ is large, the typical payment received by the agent is large. So, if the agent could have a high impact on the probability of receiving a non-zero payment, she would have incentives to work harder than the intended effort level. This would violate the first-order condition, so the probability of receiving zero payment must converge to zero (lemma A.6).
The second key step is to show that this implies that the agent’s expected utility is concave in $a$. The intuition is that the kink is very far to the left. Moreover, most of the increase in utility going from a payment of zero to the typical payment happens close to the kink. The proof shows that, due to this, the curvature of expected utility is dominated by the curvature of the cost function (lemma A.7).
6 Applications and Discussion
6.1 Closed Form Solutions
Theorem 1 implies that optimal contracts have simple functional forms in parametric settings. Contracts depend on the utility function and limited liability constraint (through the link function) and on the distribution of output (through the score). Tables 1 and 2 provide these formulae with common parameterizations. Derivations are in the supplementary material.
Table 1: Utility Functions, Link Functions, and Wage Functions
Utility Function Link Function Wage Function
$u(x)$ $g(z)$ $k\circ g(z)$
Log $\log(x+w_{0})$ $\log(\max(w_{0},z))$ $(z-w_{0})^{+}$
CRRA $\frac{(x+w_{0})^{1-\gamma}}{1-\gamma}$ $\frac{\max(w_{0}^{\gamma},z)^{\frac{1-\gamma}{\gamma}}}{1-\gamma}$ $\left((z^{+})^{\frac{1}{\gamma}}-w_{0}\right)^{+}$
CARA $\frac{-\exp(-\alpha(x+w_{0}))}{\alpha}$ $\frac{1}{\alpha\max(\exp(\alpha w_{0}),z)}$ $\frac{(\log^{+}z-\alpha w_{0})^{+}}{\alpha}$
Note: The utility function is the agent’s utility from consumption given starting wealth $w_{0}$ and a transfer $x$.
Table 2: Error Distributions
Distribution Probability Density Score Function $S(y|a)$ Mean
Gaussian $\mathcal{N}(a,\sigma^{2})$ $\frac{y-a}{\sigma^{2}}$ $a$
Log Normal $\frac{1}{y\sqrt{2\pi\sigma^{2}}}\exp\!\Bigl(-\frac{(\log(y)-a)^{2}}{2\sigma^{2}}\Bigr)$ $\frac{\log(y)-a}{\sigma^{2}}$ $\exp\!\Bigl(a+\frac{\sigma^{2}}{2}\Bigr)$
Poisson $\frac{a^{y}e^{-a}}{y!}$ $\frac{y-a}{a}$ $a$
Exponential $\frac{1}{a}e^{-\tfrac{y}{a}}$ $\frac{y-a}{a^{2}}$ $a$
Bernoulli $a^{y}(1-a)^{1-y}$, $y\in\{0,1\}$ $\frac{y-a}{a-a^{2}}$ $a$
Geometric $\Bigl(1-\frac{1}{a}\Bigr)^{y-1}\Bigl(\frac{1}{a}\Bigr)$, $y\in\{1,2,\dots\}$ $\frac{y-a}{a^{2}-a}$ $a$
Binomial $\binom{n}{y}a^{y}(1-a)^{n-y}$, $y\in\{0,\dots,n\}$ $\frac{y-na}{a-a^{2}}$ nana
Gamma $f(y\mid n,a)=\frac{y^{n-1}e^{-\tfrac{y}{a}}}{\Gamma(n)\,a^{n}}$ $\frac{y-na}{a^{2}}$ nana
Student’s $t$ $\frac{\Gamma\!\bigl(\tfrac{\nu+1}{2}\bigr)}{\Gamma\!\bigl(\tfrac{\nu}{2}\bigr)\,\sqrt{\pi\nu}\,\sigma}\left(1+\frac{1}{\nu}\,\frac{(y-a)^{2}}{\sigma^{2}}\right)^{-\tfrac{\nu+1}{2}}$ $\frac{(\nu+1)(y-a)}{\nu\,\sigma^{2}+(y-a)^{2}}$ $a$
Exponential Family $h(y)\,\exp\Bigl(\eta(a)\,T(y)-A(a)\Bigr)$ $T(y)\,\frac{d\eta(a)}{da}\;-\;\frac{dA(a)}{da}$ (Not specified)
$y=a+X,X\sim h$ g(ya)g(y-a) $\frac{g^{\prime}(y-a)}{h(y-a)}$ $a+\mathbb{E}[X]$
$y=aX,X\sim h$ $\bigl|\tfrac{1}{a}\bigr|\;h\!\Bigl(\tfrac{y}{a}\Bigr)$ $\frac{1}{a}-\frac{y}{a^{2}}\frac{g^{\prime}(\frac{y}{a})}{g(\frac{y}{a})}$ $a\,\mathbb{E}[X]$
Note: This table presents probability the PDF, score function, and means of probability distributions as functions of the agent’s chosen action, $a$.
6.2 Linear Contracts
The formulae imply piecewise linear contracts in many examples. The key ingredients are log utility, which makes the wage function piecewise linear in the score, and a linear score function. This includes an exponential family with linear sufficient statistics. That is, when $f(y,a)$ is of the form
$f(y|a)=h(y)\cdot\exp\left(\eta(a)y+A(a)\right)\text{.}$ (4)
We note this as follows:
Remark 1 (Piecewise Linear Option Contracts).
Assume that utility is log ($u(x)=\log(w_{0}+x)$ with $w_{0}>0$) and that the distribution of outcomes is in an exponential family with a linear sufficient statistic (as in equation 4). Then the canonical contract wage function is piecewise linear. This includes the Gaussian, exponential, Poisson, geometric, binomial, and gamma distributions.
Our piecewise linear contracts are option contracts. The contracts pay zero for low outcomes and increases linearly past a “strike price.” In some examples with a lower bound on the support, the optimal contract is globally linear (e.g., some of the Poisson examples in Figure 2). Finally, note that optimal contracts are only piecewise linear when the first-order approach is valid. This is illustrated in the Gaussian-log utility example in Figure 1, where the contract with negative reservation wage is clearly nonlinear.
6.3 Numerical Methods
Theorem 1 and the closed-form solutions yield efficient algorithms for solving the cost minimization problem.
Consider a relaxed cost minimization problem where we also include a finite number of global incentive compatibility constraints at a vector of actions $\boldsymbol{\hat{a}}=(\hat{a}_{1},\ldots,\hat{a}_{n})$. The Lagrangian from equation (2) becomes
[…]
$\quad+\sum\hat{\mu}_{i}\left(U(v,\hat{a}_{i})-U(v,a_{0})\right)\text{.}$
Heuristically differentiating pointwise with respect to $v(y)$ and setting the derivative to zero gives
$k^{\prime}(v(y))f(y|a_{0})=\lambda f(y|a_{0})+\mu f_{a}(y|a_{0})+\sum\hat{\mu}_{i}[f(y|a_{0})-f(y|\hat{a}_{i})]\text{.}$
The solution $v$ equals
$V(y|\lambda,\mu,\boldsymbol{\hat{\mu}},\boldsymbol{\hat{a}}):=g\biggl(\lambda+\mu S(y|a_{0})+\sum\hat{\mu}_{i}\left(1-\frac{f(y|\hat{a}_{i})}{f(y|a_{0})}\right)\biggr)\text{.}\$
Define the Lagrange dual function as
[…]
For any parametric example, the dual can be computed efficiently with the analytical formulas from section 6.1. It also has an analytic gradient by Danskin’s envelope theorem. Moreover, given a grid of outputs $\mathbf{y}_{\mathrm{grid}}$, we can cache $f(y|a_{0})$, $S(y|a_{0})$, and $1-{f(y|\hat{a}_{i})}/{f(y|a_{0})}$. This is enough to perform the numerical integrations needed for $\mathcal{D}$, so that $\mathcal{D}$ only involves matrix multiplications and applications of $g$ and $k$.
This suggests the following algorithm 1:
$\boldsymbol{\hat{a}}\leftarrow\emptyset$;  $\boldsymbol{\hat{\mu}}\leftarrow\emptyset$;  $\lambda\leftarrow\lambda_{\text{init}}$;  $\mu\leftarrow\mu_{\text{init}}$
$\mathbf{y}_{\mathrm{grid}}\leftarrow\textnormal{{InitializeGrid}}()$
repeat
 UpdateCache($\mathbf{y}_{\mathrm{grid}},\,a_{0},\,\boldsymbol{\hat{a}}$)
 // For fast dual calls
 $(\lambda,\,\mu,\,\boldsymbol{\hat{\mu}})\leftarrow\textnormal{{MaximizeDual(}}\textnormal{\emph{$\mathcal{D},\text{init}=\{\lambda,\mu,\boldsymbol{\hat{\mu}}\}$}}\textnormal{{)}}$
 // use warm start
[…]
 $\mathtt{best\_deviation},\;\mathtt{deviation\_gain}\leftarrow\textnormal{{FindBestDeviation(}}\textnormal{\emph{$v,\,a_{0}$}}\textnormal{{)}}$
 if $\mathtt{deviation\_gain}>\mathtt{tolerance}$ then
[…]
    $\boldsymbol{\hat{\mu}}\leftarrow(\boldsymbol{\hat{\mu}},0)$
    // Warm start update
   end if
until $\mathtt{deviation\_gain}\leq\mathtt{tolerance}$
return $v$
Algorithm 1 Cost minimization via active–set
Algorithm 1 solves all the examples in the paper. In our benchmarks, each run takes in the order of 3 to 14 milliseconds (see supplementary material). Performance seems to be due to three factors. First, the algorithm often ends in the first iteration because of Theorem 1. Second, dual calls are fast because of the analytic formulas and caching. Third, dual maximization is fast because of low dimensionality, analytic gradients and warm starts.
Algorithm 1 has two main drawbacks. The dual is not convex, so that the optimization step can get stuck in a local maximum. And the heuristic to add constraints might also get stuck.
To overcome these drawbacks, we implement a straightforward convex optimization solver. We discretize both the outcome and action spaces, which yields a problem with a convex objective and linear constraints. We solve this discretized problem using CVXPY with the Clarabel solver. This approach is computationally more expensive—taking between 230 and 240 milliseconds in our benchmarks—because it solves for the contract $v$ at every point of the outcome grid. However, it provides a global optimality guarantee. In our implementation, Algorithm 1 uses this convex solver as a fallback.
The main drawback of the convex solver is that the numerical optimal contract $v(y)$ is numerically unstable at points where the density $f(y|a_{0})$ is small. This is illustrated in Figure 6. The figure shows that both solvers reach essentially the same objective value. In the middle of the distribution, they reach essentially the same optimal contract. However, in the tails of the distribution, the convex solver optimal contract is numerically unstable.
Figure 6: Comparison of Algorithm 1 and convex program solvers.
Note: Top panel: optimal wage function $w(y)$. Bottom panel: agent’s expected utility $U(v^{*},a)$. Both solvers reach essentially the same expected wage cost. The convex program solution is numerically unstable in the tails where the density $f(y|a_{0})$ is small. Parameters are as in Figure 1 with $\sigma=10$. The convex solver uses a grid of 201 points for the outcomes and 200 points for the actions.
6.4 Discussion of Assumptions and Counter-Examples
Increasing score. Increasing score (also known as monotone likelihood ratio property) is a critical assumption, without which the first-order approach often fails. The importance of the assumption is easy to see with the fat-tailed Student-$t$ distribution. The $t$ distribution is similar to a Gaussian, but has fatter tails. As a consequence, the score is similar to the Gaussian score close to the mean, but approaches zero in both tails (table 2). Figure 7 displays an example where the first-order approach is invalid for any reservation wage where principal profits are positive.
Figure 7: Optimal contracts with Student-$t$ distribution and log utility.
Note: Top panel: optimal wage function $w(y)$. Bottom panel: agent’s expected utility $U(v^{*},a)$ given optimal contract and action $a$. Colors represent reservation utility. Dashed lines indicate that the first-order approach is invalid at that reservation utility. Dots indicate the recommended action. The thin horizontal line indicates indifference between local maxima. Output has Student-$t$ distribution with mean $a$, scale parameter $\sigma=20.0$, tail parameter $\nu=1.15$, initial wealth is 5050 (both in thousands of dollars), and the cost function is $c(a)=a^{2}/30000$.
Limited liability. Limited liability is crucial to guarantee existence of an optimal contract. The intuition is clear from the Lagrangian’s first order condition in equation (FOC). When the right-hand side is negative, it is optimal to give the agent the lowest possible $v(y)$. This only has a solution with both limited liability and $u(0)>$-$\infty$. This is why with log utility $u(x)=\log(w_{0}+x)$ our assumptions require $w_{0}>0$. Under our assumption that the score is unbounded below, a negative right-hand side happens for any $\mu>0$. These existence issues are well-known since [29]. Limited liability is not important for guaranteeing the validity of the first-order approach, and in fact the kink in the optimal contract makes it more difficult to guarantee the validity of the first-order approach (see proof of proposition 2).
Unbounded utility. The assumption of $u$ unbounded above is used in two places. The first is to guarantee that there exists a solution to the cost minimization problem. Otherwise, there are cases where $\bar{U}>u(\infty)-c(0)$, so that no contract satisfies the constraints of the relaxed cost minimization problem. This complicates the statements of proposition 1 (because we have to deal with the trivial infeasible case), while not adding much to the results. For example, Theorem 1 is true as stated in this case, but is vacuous, as neither problem has a solution.
The second place is showing that the probability that the agent is paid zero converges to zero as we increase reservation utility (lemma A.6). The fact that the probability converges to zero, and not just some very low positive probability, uses $u(\infty)=\infty$. In practice, however, utility functions like CARA have enough upside so that often the first-order approach is valid for many reservation wages where the problem is feasible (e.g., Figure 3).
Support of the distribution of output. Assumption 3 requires that the support of the distribution of output is the real line. This rules out interesting cases like the exponential distribution, which has a lower bound of zero. This assumption is important in the proof of Lemma A.6, that shows that the kink of the distribution approaches -$\infty$ as $\lambda$ approaches infinity. If the support of $f$ has a finite lower bound, this fails. Instead, it is possible to prove that the kink approaches the finite lower bound. The proof works in the same way, as if this did not happen it would contradict the local incentive compatibility constraint. It is an open question whether Theorem 1 is false in the case where the support of the distribution of output has a finite lower bound.
Nevertheless, we have found that the first-order approach is often valid in examples where support has a finite lower bound. For example, in Figure 2 with the Poisson distribution, the first-order approach is valid for nearly any positive reservation wage. Figure 8 displays an example with the exponential distribution, where the first-order approach is broadly valid. Likewise, we have found that the first-order approach is also broadly valid for discrete distributions, as in the Poisson example of Figure 2.
Figure 8: Optimal contracts with exponential distribution and log utility.
Note: Top panel: optimal wage function $w(y)$. Bottom panel: agent’s expected utility $U(v^{*},a)$ given optimal contract and action $a$. Colors represent reservation utility. Dashed lines indicate that the first-order approach is invalid at that reservation utility. Dots indicate the recommended action. The thin horizontal line indicates indifference between local maxima. Output has exponential distribution with mean $a$, initial wealth is 5050 (both in thousands of dollars), and the cost function is $c(a)=a^{2}/30000$.
6.5 Discussion of the Literature
First Order Approach Literature. There are three generations of contributions to the first-order approach. The first generation was started by [29] (circulated in 1975), who showed that the first-order approach does not hold generally. Prior to [29], the first-order approach was simply assumed with little justification. Subsequent papers became aware that they had to assume validity of the first-order approach [20], and developed results that do not depend on the first-order approach [17, 1].
The seminal papers in the second generation are [35] and [24]. They provided sufficient conditions for the first-order approach. Rogerson’s conditions are well known to be strict, ruling out many natural distributions, including the Gaussian. Jewitt’s conditions are more general, and include many interesting cases. Jewitt’s main assumption is implicit, in that he requires his conditions to hold for all output levels, in particular ruling out limited liability. [8] give a detailed explanation of how the Jewitt conditions are not applicable in the case of limited liability and survey the related literature. But limited liability is crucial to guarantee existence (see Section 6.4). In many examples that satisfy the Jewitt conditions, an optimal contract does not exist. This includes the Gaussian - log utility example without limited liability. At face, value, the Jewitt conditions are satisfied. But the relaxed cost minimization problem does not have a solution. Thus, the first-order approach is “valid”, since any of the non-existent solutions satisfies any property. But this validity is vacuous. This is similar to the colorful logical example that the statement “all unicorns are pink” is true, since unicorns do not exist. The second-generation papers have been influential, and much of the theoretical and applied work to this day assumes these conditions.
Our own results are inspired by [24], and we build on his insight. We impose similar conditions on the part of the contract where limited liability does not bind, while allowing for limited liability. The proof of Theorem 1 shows that, with high reservation utility, the limited liability constraint binds with low probability. We then show, in similar lines to [24], that the agent’s problem becomes concave.
There is a third generation of work extending Jewitt’s approach to multidimensional actions and richer information structures [14, 25, 6, 8, 26]. We believe our methods can be adapted to such environments, but we do not pursue that here.
At first glance, our Theorem 1 appears to contradict the standard view in the literature that the first order approach requires stringent conditions. In fact, the two are perfectly consistent. The reason is that the literature has looked for conditions under which the first-order approach is valid for every reservation utility. Our examples show that indeed any such condition must rule out basic examples like the Gaussian-log utility example from Figure 1 – and in fact all the numerical examples in this paper. Our contribution is to show that if the reservation utility is sufficiently high, the first order approach is valid.
Linear Contracts and Option Contracts. A long literature seeks conditions under which optimal contracts are simple, such as linear contracts. The classic paper is [19], who rationalize linear contracts with CARA utility in a continuous time dynamic model with Brownian output process. [22] rationalizes piecewise linear contracts with a risk neutral agent, limited liability, and a monotonicity restriction on the wage function. Notable contributions include [12] and [3] who rationalize linear contracts with risk neutral agents and a robust maxmin notion of optimality.
In contrast, we derive piecewise linear contracts under log utility and linear score. Our contracts are option contracts, with payoff of zero for low outcomes and linear increasing payment past a strike price. In some examples, our optimal contracts are linear in most of the support of the distribution of output. In some examples with a finite lower bound on the support, our optimal contracts are linear in the entire support.
The most closely related work is the contemporaneous paper [31]. He independently derived the linearity result from remark 1. [31] proves that the first-order approach is valid with an exponential distribution for some parameter values. As far as we know, [31] is the first to derive piecewise linear contracts under a broad set of conditions in standard moral hazard with risk aversion. [31] includes a complete analysis of the exponential distribution case. He characterizes optimal contracts with and without limited liability and when the optimal contract exists without limited liability. In particular, he shows that the optimal contract with limited liability is globally linear for low intended action but has a kink for high intended action. We have no overlap with these results.
In the accounting literature, we know of two papers that showed that (potentially piecewise) linear contracts are optimal under certain assumptions including the first order approach. The earliest is [18]. They assume a number of Jewitt’s conditions, and log utility with no limited liability. Their corollary 2 part 2 then shows that the optimal contract is globally linear. They prove that contracts are globally linear, whereas we prove that contracts are piecewise linear. While their result is stronger, optimal contracts often do not exist in their setting (section 6.4). The excellent graduate textbook [13] chapter 19 p. 154 also notes that piecewise linear contracts are optimal assuming that the first order approach is valid. Thus, they get to the generally correct form of piecewise linear contracts as opposed to the globally linear result. But they assume the first-order approach, and do not prove existence. Thus, [31] and our paper are the first to derive piecewise linear contracts under a broad set of conditions, rather than assuming the first-order approach. But [18] were the first to notice linearity of $k\circ g$ to the best of our knowledge. [11, 10] also attribute this observation to [18].
[…]
High Reservation Utility. In an influential paper, [5] consider the principal-agent problem in the case where reservation utility converges to infinity. This paper is one of our inspirations in considering the case of high reservation utility, although it deals with completely different issues. They assume the first-order approach is valid (their assumption 4), and give no results on validity of the first-order approach. Their main result is that the wage cost to the principal is asymptotically the same as if effort was observable, so that agency costs are asymptotically small. This is a strong result that depends on detailed assumptions that rule out many of our examples. As they note, “key assumptions driving our results are on the agent’s utility for income” … “they are not without bite: log utility is excluded, and indeed the convergence results fail in this setting.”
We now discuss the connection in detail. The gist of their results is better understood heuristically in the CRRA case, where their result requires relative risk aversion $\gamma<1$. Consider an agent with reservation wage $w_{0}$, which comprises most of the agent’s wealth. Consider first perfect information. For the principal to compensate the agent for an effort cost $c$, he must pay the agent an additional of approximately $c/u^{\prime}(w_{0})\approx cw_{0}^{\gamma}$. Therefore, as $w_{0}$ converges to infinity, the additional incentive pay is a vanishingly small fraction of the reservation wage, as long as $\gamma<1$. Their analysis formalizes how, in this case, the incentive pay is also asymptotically small with moral hazard. And this makes clear why the result depends on $\gamma<1$ and thus rules out log utility.
Their actual analysis is based on an elegant approximation. [^8] We were influenced by their approximation, but for our analysis it is more convenient to consider the average slope of $g$ over long spans of $y$ (see appendix A). The key idea is to approximate the optimal contract as a linear function of the score:
$v(y)=g\bigg(\lambda+\mu S(y|a_{0})\bigg)\approx g(\lambda)+g^{\prime}(\lambda)\mu S(y|a_{0})\text{.}$
Integrating the (IR) and (GIC) constraints yields (as in their proposition 1)
$g(\lambda)\approx c(a_{0})+\bar{U}\text{,}$
$\mu g^{\prime}(\lambda)\int S^{2}(y|a_{0})f(y|a_{0})\,dy\approx c^{\prime}(a_{0})\text{.}$
Substituting the CRRA case in the second equation implies that $\mu\lambda^{-\gamma}$ is approximately constant, and thus $\mu/\lambda$ converges to zero in the case $\gamma<1$. Their paper formalizes this point and gives the exact result. While we draw on their insight of considering high reservation utility, this discussion clarifies the differences and our marginal contribution. First, they assume the first-order approach is valid and give no results on validity of the first-order approach. Second, they make different and strong assumptions that rule out much of our analysis. Third, our results broadly hold in many settings where their asymptotics do not. For example, their results show that agency costs are asymptotically small compared to reservation wages. In all of the numerical examples in this paper, agency costs are significant, nevertheless we see that the first-order approach is broadly valid.
[4] assumes that the principal’s expected wage cost is convex in the intended action, which they motivate with a result by [5] that this is true with high stakes. But [4] does not otherwise use high reservation utility. They assume that the first-order approach is valid (their assumption 1). Neither paper assumes limited liability, and [4] assumes that utility in unbounded below but uses a bounded below utility in their numerical results. [7] consider a moral hazard problem with a set actions in an interval plus one non-ordered action. They consider the relaxed problem, assuming that the first-order approach is valid among the actions in the interval. They consider high stakes using the results in [5] to show that the optimal contract with several different utility functions converges to essentially the same contract.
Numerical Analysis. We now review existing approaches to numerically solve the moral hazard problem without the first-order approach. The review suggests that algorithm 1 offers a significant improvement. We are, to the best of our knowledge, the first to numerically solve the moral hazard problem for various realistic distributions without the first-order approach.
The literature proposes four general approaches to numerically solve the moral hazard problem without the first-order approach. The first approach is to include additional no-jump constraints in the relaxed cost minimization problem. This was introduced in the 1970s by [29] and improved by [1]. The main limitation is that there is no general method to decide which constraints to add (see [28] p. 1426). Motivated by this, [28] introduce a sandwich method where the additional constraint is added adversarially. This method works well in some special cases that are analytically tractable, such as relative risk aversion with coefficient $1/2$ where the link function is piecewise linear.
The two most popular approaches start by discretizing the set of actions, so that there is only a finite number of global incentive compatibility constraints (following [17]). The linear programming approach maximizes over a joint distribution of actions, output levels, and utility from wages, all of which have to be in a discrete grid [32, 33, 37]. While this has the advantage of making the principal’s problem a linear program, the main difficulty is the curse of dimensionality: the number of variables is the product of the sizes of the three grids. To address this, [37] introduced the mathematical program with equilibrium constraints approach. This approach maximizes over the contract and a distribution over actions. So the number of variables is the number of possible output values plus the number of possible actions.
We know of two papers who apply these approaches to realistic distributions. [33] considers a gaussian distribution with an improvement on the linear programming approach by using the Dantzig-Wolfe decomposition algorithm. [2] considers a log normal distribution, to study executive compensation, and use a mathematical program with equilibrium constraints approach. Both papers are far ahead of their time, as their numerical solutions display the key features of optimal contracts. They also report the wall time needed to solve the principal’s problem. [33] reports 51 minutes with 100 actions and 50 output levels. [2] reports that each solution takes "several hours" with 101 actions and 501 output levels. In our supplementary material benchmark table, algorithm 1 takes 93 to 356 milliseconds to solve similar Gaussian problems (without the need to discretize either actions or output levels – except for taking integrals). Naturally, these are not apples-to-apples comparisons due to improvements in hardware and implementation details. But the large performance gap suggests that algorithm 1 is substantially faster. This is consistent with algorithm 1 being significantly faster than the convex solver (see supplementary material). This is intuitive because algorithm 1 uses the analytic formulas from tables 1 and 2 and solves lower-dimensional problems.
The fourth approach in the literature is to approximate the agent’s utility as a function of actions with a ratio of polynomials [34]. [34] then use results from polynomial optimization to show that the principal’s problem can be reduced to a nonlinear program. Crucially, their method can give global optimality guarantees. They solve the log normal example from [2] in the case of a restricted contract space, with a fixed payment, equity, and options. They generously shared their code. Our benchmarks suggest that their suggested nonlinear program takes about 100 seconds in Scipy’s SLSQP solver.
Our approach is influenced by and builds upon this literature. Algorithm 1 is most related to the no-jump approach. We simply use a reasonable heuristic to iteratively add constraints, while taking advantage of Lagrange duality, parametric assumptions for fast dual calculations, analytic gradients, caching, and warm starts. Algorithm 1 has the disadvantage of not giving a global optimality guarantee, and is thus complementary to the LP and polynomial approaches. For example, one could use our algorithm to obtain a solution candidate and check it with the LP or polynomial approaches. Our convex programming approach is a natural modern extension of the linear programming and mathematical programming with equilibrium constraints approaches. Since the 2010s, there has been increasing availability of efficient convex solvers that are simple for non-experts to use, and it has become the standard tool for solving these types of problems in engineering. Our convex programming approach simply applies these new tools to the standard discretization approach from the literature going back to [17].
7 Conclusion
This paper shows that when an agent’s reservation utility is sufficiently high, the longstanding concerns about the validity of the first-order approach largely disappear. Despite the possibility of multiple local maxima under low reservation utility, forcing the principal to provide higher overall utility eliminates global deviations, ensuring both existence of an optimal contract and applicability of the standard calculus-based approach.
Appendix
Appendix A Proofs
A.1 Solution of the Relaxed Problem (Proof of Proposition 1)
Throughout this subsection, fix a reservation utility $\bar{U}$ and intended action $a_{0}$. The local IC constraint is only defined if $U(v,a)$ can be differentiated under the integral sign, so we begin by showing that differentiation under the integral at any action is possible for any feasible contract.
Remark 2.
Let $v$ be a feasible contract. Then, for all $a$ in $\mathbb{R}_{+}$, the first and second derivatives of $U(v,a)$ with respect to $a$ exist, are finite, and are given by
$U_{a}(v,a)=\int v(y)f_{a}(y|a)\,dy\text{,}$
$U_{aa}(v,a)=\int v(y)f_{aa}(y|a)\,dy\text{.}$
Proof.
The conditions for Leibniz’s rule to hold are in Assumption 2.2. ∎
We now derive some intermediate results to establish proposition 1. Recall the definition of the Lagrangian from equation (2) and definition 1 of a canonical contract.
Note that canonical contracts uniquely minimize the Lagrangian:
Lemma A.1.
Given $\lambda$ and $\mu$ in $\mathbb{R}$, there exists $v$ that minimizes the Lagrangian $\mathcal{L}(v,\lambda,\mu)$ among all feasible contracts. $v$ is $f(\cdot|a_{0})$ almost everywhere equal to the canonical contract $V(y,\lambda,\mu)$.
Proof.
The Lagrangian in equation (2) can be written as
$\mathcal{L}(v,\lambda,\mu)=\int[k(v(y))-\lambda v(y)-\mu v(y)S(y|a_{0})]f(y|a_{0})\,dy+\bar{U}\text{.}$
Differentiating the integrand with respect to $v(y)$ yields
$k^{\prime}(v(y))-\lambda-\mu S(y|a_{0})]f(y|a_{0})$
and this is strictly convex in $v(y)$. Therefore, the integrand is minimized pointwise in $v(y)$ at $v(y)=V(y|\lambda,\mu)$. Hence, the infimum is attained, and any minimizer satisfies the desired formula $f(y|a_{0})$ almost everywhere.
It only remains to show that $V(y|\lambda,\mu)$ is feasible. This follows from assumption 2 Part 4. ∎
We now note that, given $\lambda$ there is a unique value of $\mu$ that solves the local IC constraint.
Lemma A.2.
Given $\lambda$ in $\mathbb{R}$, there exists a unique $\tilde{\mu}(\lambda)$ such that the canonical contract $V(y|\lambda,\tilde{\mu}(\lambda))$ satisfies the local IC constraint (LIC). Moreover, $\tilde{\mu}(\lambda)>0$.
Proof.
We have
$U_{a}(v)=\int v(y)f_{a}(y|a_{0})\,dy\text{.}$
If $\mu=0$, then $v(y)$ is constant, so
$U_{a}(V(\cdot|\lambda,0,a_{0}))-c^{\prime}(a_{0})<0\text{.}$
As $\mu\rightarrow\infty$, $V(y|\lambda,\mu,a_{0})$ converges pointwise to $u(\infty)=\infty$ if $S(y|a_{0})>0$ and to $u(0)$ if $S(y|a_{0})<0)$. Hence, for large enough $\mu$,
$U_{a}(V(\cdot|\lambda,\mu,a_{0}))-c^{\prime}(a_{0})>0\text{.}$
Therefore, there exists at least one $\mu_{1}>0$ such that
$U_{a}(V(\cdot|\lambda,\mu_{1},a_{0}))-c^{\prime}(a_{0})=0\text{.}$
It only remains to prove that this solution $\mu_{1}$ is unique. To see this, note that
$U_{a}(V(\cdot|\lambda,\mu_{1},a_{0}))$
is weakly increasing in $\mu$. And, moreover, it is strictly increasing at any solution $\mu_{1}$ because $c^{\prime}(a_{0})>0$ implies that there is a positive measure of $y$ such that $f_{a}(y|a_{0})>0$ and
$\lambda+\mu_{1}\frac{f_{a}(y|a_{0})}{f(y|a_{0})}>\frac{1}{u^{\prime}(0)}\text{.}$
The lemma implies that the family of canonical contracts that satisfy (LIC) is a one-dimensional family indexed by $\lambda$. Define
$\tilde{V}(y|\lambda):=V(y|\lambda,\tilde{\mu}(\lambda))$ (A.1)
Define the relaxed Pareto problem as finding $v$ in $\mathcal{C}$ to
minimize $W(v)-\lambda U(v)$
subject to $\partial_{a}U(v,a_{0})=0$ (LIC)
The next lemma shows that the contracts $\tilde{V}(y|\lambda)$ span the solutions to the Pareto problem:
Lemma A.3.
Given $\lambda$ in $\mathbb{R}$, the relaxed Pareto problem has a solution, and any solution is $f(y|a_{0})$ almost everywhere equal to $\tilde{V}(y|\lambda)$.
Proof.
For any $v$ satisfying (LIC),
$W(v)-\lambda U(v)=\mathcal{L}(v,\lambda,\tilde{\mu}(\lambda))-\lambda\bar{U}\text{.}$
Lemma A.1 then implies that $W(v)-\lambda(v)$ is minimized over $\mathcal{C}$ by $\tilde{V}(y|\lambda)$, and that this solution is almost-everywhere unique. This contract satisfies (LIC) by the definition of $\tilde{\mu}(\lambda)$. ∎
Define the expected wage and utility attained by these contracts as
$\begin{array}[]{rcl}\tilde{U}(\lambda)&:=&U(\tilde{V}(\cdot|\lambda),a_{0})\text{,}\\ \tilde{W}(\lambda)&:=&W(\tilde{V}(\cdot|\lambda),a_{0})\text{.}\end{array}$
Lemma A.4.
$\tilde{U}(\lambda)$ is strictly increasing.
Proof.
Consider $\lambda_{1}<\lambda_{2}$ with optima $v_{1}:=\tilde{V}(\cdot|\lambda_{1})$ and $v_{2}:=\tilde{V}(\cdot|\lambda_{2})$. By optimality,
$W(v_{1})-\lambda_{1}U(v_{1})\leq W(v_{2})-\lambda_{1}U(v_{2})\text{,}$
and
$W(v_{2})-\lambda_{2}U(v_{2})\leq W(v_{1})-\lambda_{2}U(v_{1})\text{.}$
Adding the inequalities,
$(\lambda_{2}-\lambda_{1})\cdot\left(\tilde{U}(\lambda_{2})-\tilde{U}(\lambda_{1})\right)\geq 0\text{.}$
Therefore, $\tilde{U}$ is non-decreasing. It only remains to show that $\tilde{U}$ is strictly increasing. To reach a contradiction, assume that $\tilde{U}(\lambda_{2})=\tilde{U}(\lambda_{1})$. Optimality implies that $\tilde{W}(\lambda_{1})=\tilde{W}(\lambda_{2})$. Consider the contract $(v_{1}+v_{2})/2$. By our assumptions on the score, $v_{1}$ and $v_{2}$ differ in a set of positive measure. By strict convexity of $k$, it follows that $W((v_{1}+v_{2})/2)<W(v_{1})$. This contradicts the optimality of $v_{1}$. ∎
The proof of proposition 1 follows from collecting these results.
Proof of Proposition 1.
Let $\bar{U}_{L}:=\tilde{U}(0)$ and $\bar{U}_{R}:=\lim_{\lambda\rightarrow\infty}\tilde{U}(\lambda)$.
Part 1.
The definition of $\bar{U}_{R}$ implies that the relaxed problem is not feasible for $\bar{U}\geq\bar{U}_{R}$, as desired.
Part 2.
Let $\lambda^{*}(\bar{U})$ be 0 if $\bar{U}\leq\bar{U}_{L}$ and be the inverse of $\tilde{U}$ for $\bar{U}_{L}<\bar{U}<\bar{U}_{R}$. Lemma A.4 implies that $\lambda^{*}$ is well defined and that $\lambda^{*}(\bar{U})\geq 0$.
Let $\mu^{*}(\bar{U}):=\tilde{\mu}(\lambda^{*}(\bar{U}))$. Lemma A.2 implies that $\mu^{*}(\bar{U})>0$. Let $v^{*}(y|\bar{U}):=\tilde{V}(y|\lambda^{*}(\bar{U}))$. Note that this coincides with the definition of $v^{*}$ in the proposition statement.
Lemma A.3 implies that $v^{*}(y|\bar{U})$ solves the Pareto problem given $\lambda^{*}(\bar{U})$. This implies that $v^{*}(y|\bar{U})$ also solves the relaxed cost minimization problem given $\bar{U}$. Likewise, A.3 implies that the solution is unique almost everywhere.
Part 3, nonbinding individual rationality case.
This follows from the definition of $\lambda^{*}(\bar{U})$ which equals 0 in this range.
Part 3, binding individual rationality case. $\lambda^{*}$ strictly increasing follows from lemma A.4 and the strict convexity of $\omega$ follows from $v^{*}(\cdot|\lambda)$ being different for each value of $\lambda$ and from $k$ being strictly convex.
A.2 Concavity with High Reservation Utility (Proof of Proposition 2)
This section demonstrates that the second derivative of agent’s utility is negative for sufficiently high reservation utility. The bulk of the section is spent demonstrating the same result for $\lambda$ sufficiently high, and the main result follows as an immediate corollary. Throughout this section, we use the $\tilde{V}(y|\lambda)$ notation defined in equation (A.1). Although $\tilde{\mu}(\lambda)$ is a function of $\lambda$ (Lemma A.2) we abuse notation by omitting dependence on $\lambda$ and writing $\tilde{\mu}$ to denote $\tilde{\mu}(\lambda)$.
Formulae for derivatives of $U(v,a)$.
Lemma A.5.
Given a canonical contract, $v(y):=\tilde{V}(y|\lambda)$, the agent’s utility and its derivatives evaluated at $a$ are
$U\left(v,a\right)$ $=$ $g(\lambda)+$ $\tilde{\mu}\int\Delta g(y|\lambda)\cdot S(y|a_{0})\cdot f(y|a)\,dy-c(a),$
$U_{a}\left(v,a\right)$ $=$ $\tilde{\mu}\int\Delta g(y|\lambda)\cdot S(y|a_{0})\cdot f_{a}(y|a)\,dy-c^{\prime}(a),$
$U_{aa}\left(v,a\right)$ $=$ $\tilde{\mu}\int\Delta g(y|\lambda)\cdot S(y|a_{0})\cdot f_{aa}(y|a)\,dy-c^{\prime\prime}(a),$
where
$\Delta g(y|\lambda)=\frac{g(\lambda+\tilde{\mu}S(y|a_{0}))-g(\lambda)}{\tilde{\mu}S(y|a_{0})}.$
Proof.
The utility function, $U$, evaluated at the canonical contract, $\tilde{V}\left(y|\lambda\right)$, and action, $a$, is
$U\left(\tilde{V}(y|\lambda),a\right)=\int\tilde{V}(y|\lambda)f(y|a)\,dy.$
Substituting $\tilde{V}(y|\lambda)=g\left(\lambda+\tilde{\mu}S(y|a_{0})\right)$ yields
$U\left(\tilde{V}(y|\lambda),a\right)=\int g\left(\lambda+\tilde{\mu}S(y|a_{0})\right)f(y|a)\,dy.$
Adding and subtracting $g(\lambda)$, we rewrite the integral:
$U\left(\tilde{V}(y|\lambda),a\right)=g(\lambda)+\int\left[g\left(\lambda+\tilde{\mu}S(y|a_{0})\right)-g(\lambda)\right]f(y|a)\,dy.$
Using the definition of $\Delta g(y|\lambda)$ and multiplying by $\frac{\tilde{\mu}S(y|a_{0})}{\tilde{\mu}S(y|a_{0})}$ yields
$U\left(\tilde{V}(y|\lambda),a\right)=g(\lambda)+\tilde{\mu}\int\Delta g(y|\lambda)\cdot S(y|a_{0})f(y|a)\,dy-c(a).$
This yields the formula for $U(v,a)$. The derivatives follow from the Leibniz rule. ∎
Bounding the probability that the agent receives zero payment.
Definition A.1.
The threshold score, $\text{\@text@baccent{$S$}}(\lambda)$, is the maximum score such that the agent receives no payment. It is the score that solves
$\lambda+\tilde{\mu}\text{\@text@baccent{$S$}}(\lambda)=\frac{1}{u^{\prime}(0)}\text{,}$
or, equivalently,
$\text{\@text@baccent{$S$}}(\lambda):=\frac{1}{\tilde{\mu}u^{\prime}(0)}-\frac{\lambda}{\tilde{\mu}}\text{.}$
The threshold score is well defined because $\tilde{\mu}>0$ by Lemma A.2.
Definition A.2.
The threshold outcome, $\text{\@text@baccent{$y$}}(\lambda)$, is the outcome which induces the threshold score, $\text{\@text@baccent{$S$}}(\lambda)$:
$\text{\@text@baccent{$y$}}(\lambda):=S^{-1}(\text{\@text@baccent{$S$}}(\lambda,\tilde{\mu})|a_{0}).$
A score that satisfies the equation exists by Assumption 3.2, which states that the score’s image is $\mathbb{R}$.
Lemma A.6.
The threshold outcome approaches negative infinity as $\lambda$ approaches infinity:
$\lim_{\lambda\rightarrow\infty}\text{\@text@baccent{$y$}}(\lambda)=-\infty.$
Proof.
Lemma A.5’s result for $U_{a}$ evaluated at $a_{0}$ yields
$U_{a}\left(\tilde{V}(y|\lambda),a_{0}\right)=\tilde{\mu}\int\Delta g(y|\lambda)\cdot S(y|a_{0})f_{a}(y|a_{0})-c^{\prime}(a_{0}).$
The local incentive compatibility constraint requires that $U_{a}\left(\tilde{V}(y|\lambda),a_{0}\right)=0$. It follows that
$\tilde{\mu}\int\Delta g(y|\lambda)\cdot S(y|a_{0})f_{a}(y|a_{0})=c^{\prime}(a_{0}).$
Observe that the term inside the expectation is weakly positive because $\Delta g(y|\lambda)\geq 0$ and $S(y|a_{0})$ has the same sign as $f_{a}(y|a_{0})$. Therefore,
$c^{\prime}(a_{0})\geq\tilde{\mu}\int_{S(y|a_{0})\leq\text{\@text@baccent{$S$}}(\lambda)}\Delta g(y|\lambda)\cdot S(y|a_{0})f_{a}(y|a_{0})\,dy\text{.}$
Substituting $\Delta g(y|\lambda)$ yields
$c^{\prime}(a_{0})\geq\int_{S(y|a_{0})\leq\text{\@text@baccent{$S$}}(\lambda)}(u(0)-g(\lambda))f_{a}(y|a_{0})\,dy\text{.}$
Integrating yields:
$c^{\prime}(a_{0})\geq(u(0)-g(\lambda))\cdot F_{a}\left(\text{\@text@baccent{$y$}}(\lambda)|a_{0}\right).$
Observe that $u(0)-g(\lambda)\rightarrow$-$\infty$ by Assumption 1, and $F_{a}(\text{\@text@baccent{$y$}}(\lambda)|a_{0})<0$ by Assumption 3.2 (because the monotone likelihood ratio property implies first order stochastic dominance). Suppose $F_{a}\left(\text{\@text@baccent{$y$}}(\lambda)|a_{0}\right)$ does not approach 0. Then we have $c^{\prime}(a_{0})\geq\infty$. By contradiction, $F_{a}\left(\text{\@text@baccent{$y$}}(\lambda)|a_{0}\right)\rightarrow 0$.
The threshold outcome, $\text{\@text@baccent{$y$}}(\lambda)$, cannot converge to positive infinity because it is bounded above by the solution to $S(y|a_{0})=0$. To reach a contradiction, assume that $\text{\@text@baccent{$y$}}(\lambda)$ does not converge to negative infinity. Then there is a subsequence where $\text{\@text@baccent{$S$}}(\lambda)$ converges to a finite value $y_{0}$. Then $F_{a}\left(\text{\@text@baccent{$y$}}(\lambda)|a_{0}\right)$ converges to $F_{a}\left(y_{0}|a_{0}\right)$. Assumption 3.2 (because the monotone likelihood ratio property implies first order stochastic dominance) implies that this is a strictly negative number, reaching a contradiction. ∎
Corollary A.1.
As $\lambda\rightarrow\infty$, we have $\tilde{\mu}(\lambda)/\lambda\rightarrow 0$.
Proof.
Lemma A.6 and assumption 3.2 imply that $\text{\@text@baccent{$S$}}(\lambda)\rightarrow$-$\infty$. The result follows from the definition of $\text{\@text@baccent{$S$}}(\lambda)$. ∎
Lemma A.7.
As $\lambda$ approaches infinity, the limit of the supremum of the second derivative of agent’s utility at any action $a>0$ is strictly negative:
$\limsup_{\lambda\to\infty}\,U_{aa}\left(\tilde{V}\left(y|\lambda\right),a\right)<0\text{.}$
Proof.
By Lemma A.5,
$U_{aa}\left(\tilde{V}\left(y|\lambda\right),a\right)=g(\lambda)+\tilde{\mu}\int\Delta g(y|\lambda)\cdot S(y|a_{0})\cdot f_{aa}(y|a)\,dy-c^{\prime\prime}(a)$
Assumption 3.3 implies that there exists $y_{0}$ such that $f_{aa}(y|a)>0$ for all $y\leq y_{0}$ and $a$. Because the score is monotone, we can take $y_{0}$ such that $S(y|a_{0})<0$ for all $y\leq y_{0}$. Therefore, the integrand is negative for all $y\leq y_{0}$ and $a$.
It follows that
$U_{aa}\left(\tilde{V}\left(y|\lambda\right),a\right)\leq\tilde{\mu}\int_{y_{0}}^{\infty}\Delta g(y|\lambda)\cdot S(y|a_{0})\cdot f_{aa}(y|\hat{a})\,dy-c^{\prime\prime}(a)\text{.}$
Let $Y_{+}$ be the set of $y\geq y_{0}$ where this integrand is positive. Then
$U_{aa}\left(\tilde{V}\left(y|\lambda\right),a\right)\leq\tilde{\mu}\int_{Y_{+}}\Delta g(y|\lambda)\cdot S(y|a_{0})\cdot f_{aa}(y|\hat{a})\,dy-c^{\prime\prime}(a)\text{.}$
Lemma A.6 implies that for $\lambda$ sufficiently large, $\text{\@text@baccent{$y$}}(\lambda,a_{0})\leq y_{0}$. The concavity of $k^{\prime-1}$ (Assumption 4) implies that $\Delta g$ is decreasing for $y$ in $Y_{+}$. Therefore,
$U_{aa}\left(\tilde{V}\left(y|\lambda\right),a\right)\leq\tilde{\mu}\int_{Y_{+}}\Delta g\left(y_{0}|\lambda\right)\cdot S(y|a_{0})\cdot f_{aa}(y|\hat{a})\,dy-c^{\prime\prime}(a)\text{.}$
$g(y|\lambda)$ is concave for $y>\text{\@text@baccent{$y$}}$. Thus,
$\tilde{\mu}\Delta g(y_{0}|\lambda)$ $\leq\tilde{\mu}g^{\prime}(\lambda+\tilde{\mu}S(y_{0}|a_{0}))$
$=\frac{\tilde{\mu}}{\lambda+\tilde{\mu}S(y_{0}|a_{0})}\cdot(\lambda+\tilde{\mu}S(y_{0}|a_{0}))g^{\prime}(\lambda+\tilde{\mu}S(y_{0}|a_{0})).$
Corollary A.1 states that $\tilde{\mu}/\lambda\rightarrow 0$ as $\lambda\rightarrow\infty$, and Assumption 4 implies that $(\lambda+\tilde{\mu}S(y_{0}|a_{0}))g^{\prime}(\lambda+\tilde{\mu}S(y_{0}|a_{0}))$ has a finite limit. It follows that
$\limsup_{\lambda\to\infty}\,U_{aa}\left(\tilde{V}\left(y|\lambda\right),a\right)\leq-c^{\prime\prime}(a)<0$
with the last inequality implied by the strict convexity of the cost function (Assumption 1). ∎
Completing the proof of Proposition 2.
Proof of Proposition 2.
We first prove that there exists $\lambda_{0}$ such that, for all $\lambda\geq\lambda_{0}$ and $a$ in $\mathcal{A}$,
$U_{aa}\left(\tilde{V}(y|\lambda),a\right)\leq 0\text{.}$
To reach a contradiction, assume that this is not the case. Then there exists a sequence of $\lambda_{n}\rightarrow\infty$ and $a_{k}$ in $\mathcal{A}$ such that
$U_{aa}\left(\tilde{V}(y|\lambda_{n}),_{n}\right)>0\text{.}$
Because $\mathcal{A}$ is compact, we can take a convergent subsequence where $a_{k}\rightarrow a_{1}$. Therefore,
$\limsup_{\lambda\rightarrow\infty}U_{aa}(\tilde{V}(y|\lambda),a_{1})>0\text{.}$
This contradicts Lemma A.7.
For any $\lambda\geq\lambda_{0}$, we thus have that $U(\tilde{V}(y|\lambda),a)$ is concave in $a$. The proposition is proven by letting $U^{*}$ be the solution to $\lambda^{*}(U^{*})=\lambda_{0}$. The agent’s problem is concave in $a$ for $\bar{U}\geq U^{*}$ because $\lambda^{*}(\bar{U})$ is monotonic by Proposition 1. ∎